Theresa sold three different magazine subscribtions for $28,$18,and $27. She sold three fewer of the $28 subscriptions than of the $18 subscriptions and a total of 34 subscriptions. If her total sales wer $819, how many $27 subscriptions did Theresa sell?
Answer by stanbon(48568)

Theresa sold three different magazine subscribtions for $28,$18,and $27. She sold three fewer of the $28 subscriptions than of the $18 subscriptions and a total of 34 subscriptions. If her total sales wer $819, how many $27 subscriptions did Theresa sell?
-----------
Let number of of $18 subs be "x"; value is 18x dollars
Number of $28 subs is "x-3"; value is 28(x-3)=28x-84 dollars
Number of $27 subs is "34-x-(x-3): value is 27(37-2x)=999-54x dollars
-------------------
EQUATION:
value + value + value = $819
18x+28x-84+999-54x=819
-8x=-96
x=12 subsciptions sold at $18
37-2x==37-24=13 substitutions sold at $27
------------------
Cheers,
Stan H.